OpenStudy (anonymous):

An angle measure of 82 degrees is equivalent to ____ radians. Round your answer to the nearest hundredth when necessary.

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

when you need to convert degree to radius, just multiple degree by \(\dfrac{\pi}{180^o}\).

OpenStudy (anonymous):

so how would i do this one

geerky42 (geerky42):

Like this: \[82^o \times \dfrac{\pi}{180^o}\]

OpenStudy (anonymous):

1.43044444444 this is what i got

geerky42 (geerky42):

yep, now round your answer to the nearest hundredth, then you are done.

OpenStudy (anonymous):

how do i do that just round it up to 1.45

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Actually, 1.43044444444... rounds up to 1.43 because digit after 3 is 0.